Could you please show work on how you get the answer for: One number is 4 less than 3 times another. Their sum is -16. Find the smaller.
Answer by oberobic(2304)

Let x and y be two numbers.
.
We're told x is 4 less than three times y.
.
x = 3y-4
.
We're told the sum is -16.
.
x + y = -16
so
x = -16-y
.
Substitute x = 3y-4
.
3y-4 = -16-y
4y = -12
y = -3
.
Substitute to find 'x'
.
x + y = -16
x + (-3) = -16
x = -13
.
Check the solution to be sure it is the answer.
.
Does x = 3y-4?
.
3(-3)-4 = -9-4 = -13
Correct.
.
Answer: The smaller number is -3.
.
Done.